OpenStudy (anonymous):

7t^6*8t^4 were do I start? Since the terms are NOT like terms??

OpenStudy (anonymous):

remember a^b*a^c=a^(b+c) hence 7*8(t^6*t^4) = 56(t^(6+4)) =56(t^10)

OpenStudy (anonymous):

(7*8)t^(6+4) =56t^10

OpenStudy (anonymous):

got it jess-ca

OpenStudy (anonymous):

THANKS :) Yup got it

OpenStudy (anonymous):

=)